express 31 as the sum of three odd primes.
Found 3 solutions by amarjeeth123, josgarithmetic, Alan3354:
Answer by amarjeeth123(569)

31 can be written as 3+11+17.
The three odd primes are 3,11 and 17.

Answer by josgarithmetic(39617)
Answer by Alan3354(69443)

express 31 as the sum of three odd primes.
---------
5+7+19 = 31
5+13+13 = 31
7+7+17 = 31
7+11+13 = 31